OpenStudy (yolomcswagginsggg):

(1 pt) Which related equation can be used to solve the equation? Column A Column B 1. –8.7 = k + 6 2. –6 = k + 8.7 3. 8.7 = k + (–6) 4. 6 = k + 8.7 A. k = –8.7 – 6 B. k = 6 – 8.7 C. k = 8.7 – (–6)

OpenStudy (yolomcswagginsggg):

colum b + A. k = –8.7 – 6 B. k = 6 – 8.7 C. k = 8.7 – (–6)